Circular No. 8912

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 A2 (LINEAR) An apparently asteroidal object discovered by the LINEAR survey

(discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to show cometary appearance on images obtained at Table Mountain: CCD images on Jan. 16.24–16.26 UT by J. Young (0.61-m reflector) show a 5′′-6′′ round coma of mag 19.0 and a bright central condensation in marginal seeing, and a 4′′ coma is visible on images from Jan. 19.1. Young adds that the comet has changed brightness considerably, being at mag 19.5 on Jan. 14, 19.3 on Jan. 18, and 20.2 on Jan. 19 — though the coma is less pronounced on these other dates (the Jan. 18 images were obtained by D. Mayes).

2008 UT  α_{2000}  δ_{2000}  Mag.
Jan. 13.22223  6^{h}03^{m}22.^{s}94  +43^{o}46′02′′.8  20.1

The available astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2008-B18.

T = 2008 June 12.932 TT  ω = 233.^{o}546
e = 0.59071  Ω = 315.368 }2000.0
q = 1.30603 AU  i = 19.260

a = 3.19100 AU n^{o} = 0.172908 P = 5.70 years

COMETS C/2007 T12–T14, C/2007 U3, C/2007 U4 (SOHO) Further to IAUC 8910, additional Kreutz-sungrazing comets have been

found on SOHO website images, as tabulated below. C/2007 T12 was small and diffuse (mag 7.5). C/2007 T13 was extremely faint (mag ∼ 8) and diffuse. C/2007 T14 was extremely faint (mag ∼ 7.5) and tiny in C3coronagraph images, and stellar in appearance (mag 7.5) in C2 images. C/2007 U3 was stellar in appearance (mag ∼ 6) in C3 images, and fairly condensed with a hint of a thin, faint tail in C2 images. C/2007 U4 was diffuse and elongated (mag 7.5).

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 T12 Oct. 13.579 13^{h}06.^{m}4 − 8^{o}43′ C2  AK 2008-A50
C/2007 T13 15.329 13 12.4 − 9 11 C2 HS 2008-A50
C/2007 T14  15.471 13 14.9 − 9 55 C3/2 HS 2008-A50
C/2007 U3  24.113 13 37.3 −13 35 C3/2 HS 2008-A51
C/2007 U4  24.146 13 46.9 −12 39 C2  HS 2008-A51
2008 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green